Qweak target cool-down had a setup around lunch time: cool-down 4K transfer line caused instability at ESR which took down 15K supply. Their target warmed up to 40K, 50 psi. Re-cool-down started around 2 pm with mainly 4K, then 15K kicked in after the inlet became cold enough. It is still on going with no good estimation when it would finish. We need to wait for the Qweak target cool-down complete and the ESR becomes stable again. Neither Hall C nor cryo is willing to give me an estimation when this will be (late tonight, tomorrow,...). Stay tuned.